OpenStudy (anonymous):

An object is thrown directly up (positive direction) with a velocity (vo) of 20.0 m/s and do= 0. How high does it rise (v = 0 cm/s at top of rise). Remember, acceleration is -9.80 m/s2. d = _____ m

OpenStudy (anonymous):

By v^2 = u^2 + 2as U =20 V=0 -9.81 =a S = 20.387 m